[Brahms-l] [Fwd: Shift report 20031221 00:00-08:00]

From: Flemming Videbaek <videbaek@rcf.rhic.bnl.gov>
Date: Sun Dec 21 2003 - 16:31:03 EST
-------- Original Message --------
Subject: Shift report 20031221 00:00-08:00
From: Apache <apache@pii3.brahms.bnl.gov>
Date: Sun, December 21, 2003 6:33 am
To: brahms-l@lists.bnl.gov, videbaek@bnl.gov


------------------------------------------------------------
Date: Sun Dec 21 06:33:36 2003, Shift: 00:00-08:00
Supervisor: hiro, with: catalin and oana
Log book#15 pages 12 - 12

Shift Summary:
------------------------------------------------------------
1.  beam from 1am. (1 store)
2.  run 9217 should be used to calibrate bb and zdc if TPM1 is working.
(It looks like it is working but its drift voltage is at 4450V instead
of 4900V)  T3 does not hold voltage despite the several attempts.
TPM2 is not on since it has pulser setting.
c1/c4/h1/h2 high voltage does not communicate.
Attenuation setting in PTQ for silicon has been adjusted.
3. E-Log spell checker function is finally working. (I need to put spell
checker in shift report page too.)
   shift report should now be sent to our mailing list.  (let's see if
this shift report shows up in our list.)

------------------------------------------------------------


-- 
Flemming Videbaek
Brookhaven National Laboratory
Physics Department
tlf 631-344-4106
fax 631-344-1334
e-mail: videbaek@bnl.gov




_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Dec 21 16:32:06 2003

This archive was generated by hypermail 2.1.8 : Sun Dec 21 2003 - 16:32:44 EST